# In Search of the Long-Tail: Systematic Generation of Long-Tail Inferential Knowledge via Logical Rule Guided Search

### v2

@@ 6 Conclusion and Future Work
Appendix A Symbolic Rule Creation
Following the criteria mentioned in § 2.2, we curated 417 symbolic rules from the following sources:
@@ Appendix A Symbolic Rule Creation
    Expansion from an abstract rule. We prompt GPT4 to obtain concrete data types for the abstract rule to create multiple concrete rules. Table 5 shows an example of one abstract rule, related prompts for expansion and one concrete rule. We obtained 201 rules using this method.
@@
    Direct generation from concrete data type pairs. We prompt GPT4 to generate multiple predicates defining the relationship of the data types as the conclusion, and then generate a premise. We obtained 216 rules using this method.
Table 5: An example meta-rule and its corresponding prompts for expansion.
Meta-Rule
is_of_age(Person A, Age X) & requires_a_ minimal_age_of(Object B, Age Y) & is_smaller_than(Age X, Age Y) $\to$ cannot_operate(Person A, Object B)
Prompt for Data Type Expansion
In rule “requires_a_minimal_age_of(Object B, Age Y) & cannot_operate(Person A, Object B)", B is a variable representing an object. List 10 subcategories of object that B could be that also make the rule true.
Expanded Data Types
Vehicle, Machinery, Alcohol, Firearm, Tattoo Equipment, Tobacco Product
Prompt for Verb Optimization
cannot_operate(Person A, Object B) is equal to [mask](Person A, Vehicle B). Write the best predicate that could fit in [mask] token.
Expanded Rule
is_of_age(Person A, Age X) & requires_a_ minimal_operating_age_of(Object B, Age Y) & is_smaller_than(Age X, Age Y) $\to$ cannot_drive(Person A, Vehicle B)
Appendix B Critic Model
@@
Appendix C Ablation Studies on LINK
C.1 Hyperparameters on Knowledge Model
@@
C.2 Effect of Critic on LINK
To investigate the effectiveness of the critic, we provide an ablation study on a few sampled rules by removing the critic in LINK. Table 8 shows the generation quality of LINK and several variants in long-tail distribution. Without the critic, the generation quality decreases significantly. However, the performance drop is less significant in the head distribution Table 9. Besides, if we replace the reranker with a random sampling method, the generated statements cannot lie in the long-tail distribution (which will be further explained in § C.3) and have higher correctness without the critic. It indicates that it is harder for models to generate correct statements from the long-tail distribution than the head distribution without LINK.
@@ C.2 Effect of Critic on LINK
C.3 Effect of Reranker on LINK
@@
C.4 Ineffectiveness of Post-hoc Reranking for LLM generated knowledge.
@@
C.5 Applying GPT4 as the knowledge model
@@
Appendix D Addendum on Distribution
@@
D.1 Additional distribution plots for symbolic rules
@@
D.2 Distribution Comparison of Different Models
@@
Appendix E Entailment Classification Probing
E.1 Probing Template
@@ 7 Answer the question with yes or no: if premise, conclusion_question? Depends
E.2 Rationale analysis
@@
Input
Premise: PersonX is of age 17 years 8 months. Conclusion: PersonX cannot consume Bourbon. Does premise entail conclusion?
Label
Yes
Answer
No
Rationale
In many countries, the legal drinking age is 18 or 21. If we do not know the specific country PersonX is in, we cannot definitively conclude that they cannot consume Bourbon.
Appendix F Domain-wise statistics of LINT
@@
F.1 Domain-wise human evaluation
@@ F.2 Rule definitions
Rule0
lived_in(Person P, Geographic Location A) & lived_during(Person P, Historical Time Period D) & existed_during(Geographic Location A, Historical Time Period D) & was_invented_in(Product or Technology C, Year Y) & is_more_than_a_century_earlier_ than(Historical Time Period D, Year Y) $\to$ is_not_able_to_use(Person P, Product or Technology C)
Rule32
has_trouble_lifting(Person X, Name of Appliance B) & is_heavier_than(Object A, Name of Appliance B) $\to$ cannot_lift(Person X, Object A)
Rule46
is_allergic_to(Person A, Substance X) & includes(Name of Cosmetics B, Substance X) $\to$ cannot_use(Person A, Name of Cosmetics B)
Rule88
died_in(Historical Figure A, Historical Time Period X) & was_created_during(Artifact B, Historical Time Period Y) & is_earlier_than( Historical Time Period X, Historical Time Period Y) $\to$ cannot_create(Historical Figure A, Artifact B)
Rule112
has_trouble_containing(Room B, Furniture C) & is_larger_than(Furniture A, Furniture C) $\to$ cannot_fit_in(Furniture A, Room B)
Rule122
has_trouble_containing(Trunk B, Furniture C) & is_larger_than(Furniture A, Furniture C) $\to$ cannot_fit_in(Furniture A, Trunk B)
Appendix G Amazon Mechanic Turk
G.1 Recruiting Workers
We recruit workers from all English-speaking countries (US, UK, New Zealand, Australia, Canada), although AMT workers are mostly US-based. We use a qualification task to recruit AMT workers. In the qualification task, all workers will be presented with three manually selected statements, which are clear and representative. Each statement has five related questions as described in Appendix B. Only workers who answer all the questions correctly will be recruited. In the end, we recruited 38 workers to evaluate the quality of generation and 17 workers as human baselines for the entailment classification task. We paid the workers \$0.47 per annotation for evaluating the quality of generations and \$0.11 per annotation for the entailment classification task, to match \$15 per hour based on their working time.
G.2 Templates
@@ G.3 Agreement statistics
G.4 Failure Case Examples
@@
Rule 172
Locational
Rule: is_located_in(Person A, Location X) & is_forbidden_in(Food Item B, Location X) $\to$ cannot_eat(Person A, Food Item B) Premise: Person X is located in Houston Conclusion: Person X cannot eat Chocolate Is Houston a location? Annotation: Yes Is Chocolate a food item? Annotation: Yes Does the premise entail the conclusion? Annotation: No Reason: It is a factual error. Chocolate is not actually forbidden in Houston, so People in Houston can eat chocolate.
Rule 371
Capability and Advice
Rule: can_treat(Drug B, Name of Disease X) & has(Person A, Name of Disease X) $\to$ should_take(Person A, Drug B) Premise: Person X has Hepatitis Conclusion: Person X should take Sofosbuvir Is Hepatitis a name of disease? Annotation: Yes Is Sofosbuvir a drug? Annotation: Yes Does the premise entail the conclusion? Annotation: No Reason: It is a factual error. There are different types of hepatitis viruses. Sofosbuvir is a medication used primarily for the treatment of hepatitis C. For other types of hepatitis, different medications or treatments may be necessary.
Rule 274
Temporal
Rule: vanished_in(Plant A, Historical Time Period X) & was_invented_in(Weapon B, Historical Time Period Y) & is_earlier_than(Historical Time Period X, Historical Time Period Y) $\to$ cannot_be_used_to_conceal(Plant A, Weapon B) Premise: Plant X vanished in Mongol Conclusion: Plant X cannot be used to conceal M92 Zolja Is Mongol a historical time period? Annotation: No Is M92 Zolja a weapon? Annotation: Yes Does the premise entail the conclusion? Annotation: Yes Reason: It is a data type error. The Mongols are an East Asian ethnic group native to Mongolia, not a time period. The Mongol Empire may refer to a period of the 13th and 14th centuries, but Mongol cannot.
Rule 204
Natural Properties
Rule: has_trouble_containing(Drawer B, Tool C) & is_larger_than (Tool A, Tool C) $\to$ cannot_be_placed_in(Tool A, Drawer B) Premise: Drawer X has trouble containing Scroll saw Conclusion: Car cannot be placed in Drawer X Is Scroll saw a Tool? Annotation: Yes Is Car a Tool? Annotation: No Does the premise entail the conclusion? Annotation: Yes Reason: It is a data type error. Car is a vehicle instead of a tool.
Refer to caption Figure 13: AMT template for the evaluation of generation quality. Refer to caption Figure 14: AMT template for the human baseline of the entailment classification task.

### v3

@@ 6 Conclusion
Appendix A More Discussions
Our main experiments on entailment classification task assume that the LLMs have the specific knowledge necessary to answer the questions. While this is a fair implication, given that their training data are known to contain essentially anything available on the web, one may also worry that the model’s handling of such knowledge is impacted by the memorization of this knowledge. As knowledge gets more long-tail , it means it appears in the training data less and is thus harder to memorize. Given the imperfect memorization of current LLMs, this may impact their performance on such knowledge, and our current experiments suggest that. To verify this, for future experiments one can count the frequency of this long-tail knowledge in the training data to measure how imperfect memorization makes certain knowledge long-tail .
Appendix B Symbolic Rule Creation
Following the criteria mentioned in § 2.2, we curated 417 symbolic rules using the following steps:
@@ Appendix B Symbolic Rule Creation
    Select Compatibility (conclusion will be positive) or Mutual Exclusivity (conclusion will be negative).
@@
    Defining Constraints. Our constraints can be categorized as temporal, locational, natural properties, and desirable/undesirable outcomes and effects. Temporal constraints refer to time period or age, locational constraints refer to geographic location such as countries or cities as well as climates (tropical or polar, etc), natural properties refer to physical properties of objects such as temperature, size, density, speed, and outcomes and effects include allergies, cure of disease, etc.
  3. 3.
    Selecting argument types. We either use Person-Object or Object-Object as key arguments.
  4. 4.
    Defining interactions between arguments. Person-Object interactions include “using”, “operating”, “buying”, “consuming”, and Object-Object are more type-specific (eg. “scratching” when the constraint is “hardness” in natural property).
  5. 5.
    Expanding data types for Person or Object. We prompt InstructGPT to generate more specific data types under Person (eg. Historical Figure) or Object (eg. Vehicle, Tool).
  6. 6.
    Optimize the verb describing the interaction. We prompt InstructGPT to generate a more accurate verb for the expanded data types.
Table 5 shows how an example symbolic rule is constructed.
Table 5: Illustration of our process for creating an example symbolic rule.
Constraint Temporal (Age)
Arguments Person-Object
Constraint Predicate is_of_age(Person A, Age X) & requires_a_ minimal_age_of(Object B, Age Y) & is_smaller_than(Age X, Age Y)
Principle Mutual Exclusivity
Interaction cannot_operate(Person A, Object B)
Prompt for Data Type Expansion In rule “requires_a_minimal_age_of(Object B, Age Y) & cannot_operate(Person A, Object B)", B is a variable representing an object. List 10 subcategories of object that B could be that also make the rule true.
Expanded Data Types Vehicle, Machinery, Alcohol, Firearm, Tattoo Equipment, Tobacco Product
Prompt for Verb Optimization cannot_operate(Person A, Object B) is equal to [mask](Person A, Vehicle B). Write the best predicate that could fit in [mask] token.
Expanded Rule Example is_of_age(Person A, Age X) & requires_a_ minimal_operating_age_of(Object B, Age Y) & is_smaller_than(Age X, Age Y) $\to$ cannot_drive(Person A, Vehicle B)
Appendix C Critic Model
@@
Appendix D Ablation Studies on LINK
D.1 Hyperparameters on Knowledge Model
@@
D.2 Effect of Critic on LINK
To investigate the effectiveness of the critic, we provide an ablation study on a few sampled rules by removing the critic in LINK. Table 8 shows the generation quality of LINK and several variants in long-tail distribution. Without the critic, the generation quality decreases significantly. However, the performance drop is less significant in the head distribution Table 9. Besides, if we replace the reranker with a random sampling method, the generated statements cannot lie in the long-tail distribution (which will be further explained in § D.3) and have higher correctness without the critic. It indicates that it is harder for models to generate correct statements from the long-tail distribution than the head distribution without LINK.
@@ D.2 Effect of Critic on LINK
D.3 Effect of Reranker on LINK
@@
D.4 Ineffectiveness of Post-hoc Reranking for LLM generated knowledge.
@@
D.5 Applying GPT4 as the knowledge model
@@
Appendix E Addendum on Distribution
@@
E.1 Additional distribution plots for symbolic rules
@@
E.2 Distribution Comparison of Different Models
@@
Appendix F Long-tail Prompts for LLMs
We tried 10 prompts when prompting LLMs to generate knowledge statements in the long-tail distribution directly with instructions. Table 11 shows the 10 prompts which are appended to the original instruction.
Table 11: An illustration of 10 prompts that instruct LLMs to generate knowledge statements in the long-tail distribution.
Prompt
1 Use less frequent terms of A and B and Z.
2 Use terms of A and B and Z that are less common.
3 Use terms with lower frequency for A and B and Z.
4 Use terms of A and B and Z that have lower probability in language model distribution.
5 Use less frequent words of A and B and Z.
6 Use words of A and B and Z that are less common.
7 Use words with lower frequency for A and B and Z.
8 Use less frequent entities of A and B and Z.
9 Use entities of A and B and Z that are less common.
10 Use entities with lower frequency for A and B and Z.
Figure 13: LINT has a higher $\delta$ than GPT4 baselines with 10 different prompts.
Following Section 3.2, Figure 13 shows the $\delta$ value for GPT4 baselines with these prompts and LINT. These prompts have a similar effect on the distribution of generated statements on most rules and LINT consistently has a higher $\delta$, indicating that no matter the prompt we use, LLMs cannot directly generate long-tail statements by following instructions. Though these prompts have different $\delta$ values on some rules, no one can consistently achieve a higher $\delta$ among them.
Appendix G Entailment Classification Probing
G.1 Probing template
@@ 7 Answer the question with yes or no: if premise, conclusion_question? Depends
G.2 Accuracy averaged over templates
One main concern of not using accuracy of our main metric is because accuracy is not favorable for label imbalance. Because we have 13 templates where for each knowledge statement we will have an uneven number of positive or negative answers, the negative label rate in total is around 51.58%. Therefore, we decide to consider the template biases and deem a model correct only if it answers all templates of a statement correctly, as our goal is to test the true knowledge of models.
When evaluating the accuracy across all templates, the model’s performance also drops over long-tail knowledge. Table 13 shows the accuracy across all templates of GPT4. Even if LLMs are brittle to templates, they exhibit a performance drop in the long-tail distribution among all domains.
Head Long-tail
Natural Properties 67.07 63.40
Temporal 83.14 78.97
Outcomes and Effects 81.79 80.19
Locational 68.90 67.07
Table 13: Accuracy over all templates, GPT4
G.3 Model and human error analysis
Table 14 shows probing examples in which models and humans make error. For those from the head distribution, the entailment can be easy if humans can use search engines. For example, humans can search the periods of “The Paleolithic Era” and “Lion Gate of Mycenae”, and answer easily. Thus the human errors in the head distribution may be due to carelessness. For those containing long-tail knowledge, even with search engines, it is not so easy to infer the answer for human annotators. It is also likely that models do not have such knowledge either.
Table 14: Examples that human and model make mistakes.
Statement: Yes or no: if Plant X vanished in Paleolithic Era, Plant X cannot surround Lion Gate of Mycenae. (Head)
True label: Yes
Human answer: No
Model answer (GPT4 + COT): Yes. The Paleolithic Era ended around 10,000 BC and the Lion Gate of Mycenae was built around 1250 BC. If PlantX vanished in the Paleolithic Era, it could not have surrounded the Lion Gate of Mycenae which was built much later.
Statement: Yes or no: if Bag X has trouble containing Pandeiro, Dhak cannot fit in Bag X. (Long-tail )
True label: Yes
Human answer: No
Model answer (GPT4 + COT): No. The size of the Pandeiro does not directly relate to the size of the Dhak. They are different objects and may have different shapes and sizes. Therefore, just because BagX has trouble containing a Pandeiro doesn’t necessarily mean that a Dhak cannot fit in BagX.
G.4 Rationale analysis
@@
Input Premise: PersonX is of age 17 years 8 months. Conclusion: PersonX cannot consume Bourbon. Does premise entail conclusion?
Label Yes
Answer No
Rationale In many countries, the legal drinking age is 18 or 21. If we do not know the specific country PersonX is in, we cannot definitively conclude that they cannot consume Bourbon.
Appendix H Domain-wise statistics of LINT
@@
H.1 Domain-wise human evaluation
@@ H.2 Rule definitions
Rule0 lived_in(Person P, Geographic Location A) & lived_during(Person P, Historical Time Period D) & existed_during(Geographic Location A, Historical Time Period D) & was_invented_in(Product or Technology C, Year Y) & is_more_than_a_century_earlier_ than(Historical Time Period D, Year Y) $\to$ is_not_able_to_use(Person P, Product or Technology C)
Rule32 has_trouble_lifting(Person X, Name of Appliance B) & is_heavier_than(Object A, Name of Appliance B) $\to$ cannot_lift(Person X, Object A)
Rule46 is_allergic_to(Person A, Substance X) & includes(Name of Cosmetics B, Substance X) $\to$ cannot_use(Person A, Name of Cosmetics B)
Rule88 died_in(Historical Figure A, Historical Time Period X) & was_created_during(Artifact B, Historical Time Period Y) & is_earlier_than( Historical Time Period X, Historical Time Period Y) $\to$ cannot_create(Historical Figure A, Artifact B)
Rule112 has_trouble_containing(Room B, Furniture C) & is_larger_than(Furniture A, Furniture C) $\to$ cannot_fit_in(Furniture A, Room B)
Rule122 has_trouble_containing(Trunk B, Furniture C) & is_larger_than(Furniture A, Furniture C) $\to$ cannot_fit_in(Furniture A, Trunk B)
Appendix I Amazon Mechanic Turk
I.1 Recruiting Workers
We recruit workers from all English-speaking countries (US, UK, New Zealand, Australia, Canada), although AMT workers are mostly US-based. We use a qualification task to recruit AMT workers. In the qualification task, all workers will be presented with three manually selected statements, which are clear and representative. Each statement has five related questions as described in Appendix C. Only workers who answer all the questions correctly will be recruited. In the end, we recruited 38 workers to evaluate the quality of generation and 17 workers as human baselines for the entailment classification task. We paid the workers \$0.47 per annotation for evaluating the quality of generations and \$0.11 per annotation for the entailment classification task, to match \$15 per hour based on their working time.
I.2 Templates
@@ I.3 Agreement statistics
I.4 Failure Case Examples
@@
Rule 172 Locational Rule: is_located_in(Person A, Location X) & is_forbidden_in(Food Item B, Location X) $\to$ cannot_eat(Person A, Food Item B) Premise: Person X is located in Houston Conclusion: Person X cannot eat Chocolate Is Houston a location? Annotation: Yes Is Chocolate a food item? Annotation: Yes Does the premise entail the conclusion? Annotation: No Reason: It is a factual error. Chocolate is not actually forbidden in Houston, so People in Houston can eat chocolate.
Rule 371 Capability and Advice Rule: can_treat(Drug B, Name of Disease X) & has(Person A, Name of Disease X) $\to$ should_take(Person A, Drug B) Premise: Person X has Hepatitis Conclusion: Person X should take Sofosbuvir Is Hepatitis a name of disease? Annotation: Yes Is Sofosbuvir a drug? Annotation: Yes Does the premise entail the conclusion? Annotation: No Reason: It is a factual error. There are different types of hepatitis viruses. Sofosbuvir is a medication used primarily for the treatment of hepatitis C. For other types of hepatitis, different medications or treatments may be necessary.
Rule 274 Temporal Rule: vanished_in(Plant A, Historical Time Period X) & was_invented_in(Weapon B, Historical Time Period Y) & is_earlier_than(Historical Time Period X, Historical Time Period Y) $\to$ cannot_be_used_to_conceal(Plant A, Weapon B) Premise: Plant X vanished in Mongol Conclusion: Plant X cannot be used to conceal M92 Zolja Is Mongol a historical time period? Annotation: No Is M92 Zolja a weapon? Annotation: Yes Does the premise entail the conclusion? Annotation: Yes Reason: It is a data type error. The Mongols are an East Asian ethnic group native to Mongolia, not a time period. The Mongol Empire may refer to a period of the 13th and 14th centuries, but Mongol cannot.
Rule 204 Natural Properties Rule: has_trouble_containing(Drawer B, Tool C) & is_larger_than (Tool A, Tool C) $\to$ cannot_be_placed_in(Tool A, Drawer B) Premise: Drawer X has trouble containing Scroll saw Conclusion: Car cannot be placed in Drawer X Is Scroll saw a Tool? Annotation: Yes Is Car a Tool? Annotation: No Does the premise entail the conclusion? Annotation: Yes Reason: It is a data type error. Car is a vehicle instead of a tool.
Figure 14: AMT template for the evaluation of generation quality.
Figure 15: AMT template for the human baseline of the entailment classification task.